Reduce the equation x²(y-px)=yp² to clairaut´s form and hence find its complete solution.
1
Expert's answer
2019-08-19T13:35:42-0400

Let's put "X=x^2;\quad Y=y^2" then "p=\frac{x}{y}\frac{dY}{dX}" . Let's denote "P=\frac{dY}{dX}"

Then the equation can be rewritten in form

"X(y-x^2\frac{P}{y})=y\frac{x^2}{y^2}P^2"

By multiplying both sides with y, we assume

"X(y^2-x^2P)=x^2P^2"

Or

"X(Y-XP)=XP^2"

Therefore

"Y=XP+P^2"

which is now in Clairaut’s form

The solution got by just replacing P by constant c.

Hence

"Y=cX+c^2"

or

"y^2=cx^2+c^2"